40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 22 02:03:57 2010, rana, Summary, Electronics, UPDH Box #17: Ready 
    Reply  Mon Mar 22 10:07:30 2010, Koji, Summary, Electronics, UPDH Box #17: Ready 15V_power_supply.pdf
Message ID: 2692     Entry time: Mon Mar 22 02:03:57 2010     Reply to this: 2693
Author: rana 
Type: Summary 
Category: Electronics 
Subject: UPDH Box #17: Ready 

It took too long to get this box ready for action. I implemented all of the changes that I made on the previous one (#1437). In addition, since this one is to be used for phase locking, I also made it have a ~flat transfer function. With the Boost ON, the TF magnitude will go up like 1/f below ~1 kHz.

The main trouble that I had was with the -12V regulator. The output noise level was ~500 nV/rHz, but there was a large oscillation at its output at ~65 kHz. This was showing up in the output noise spectrum of U1 (the first op-amp after the mixer). Since the PSRR of the OP27 is only ~40 dB at such a high frequency, it is not strange to see the power supply noise showing up (the input referred noise of the OP27 is 3.5 nV/rHz, so any PS noise above ~350 nV/rHz becomes relavent).

I was able to tame this by putting a 10 uF tantalum cap on the output of the regulator. However, when I replaced the regulator with a LM7912 from the blue box, it showed an output noise that went up like 1/f below 50 kHz !! I replaced it a couple more times with no benefit. It seems that something on the board must now be damaged. I checked another of the UPDH boxes, and it has the same high frequency oscillation but not so much excess voltage noise. I found that removing the protection diode on the output of the regulator decreased the noise by a factor of ~2. I also tried replacing all of the 1 uF caps that are around the regulator. No luck.

Both of the +12 V regulators seem fine: normal noise levels of ~200 nV/rHz and no oscillations.

Its clear that the regulator is not functioning well and my only guess is that its a layout issue on the board or else there's a busted component somewhere that I can't find. In any case, it seems to be functioning now and can be used for the phase locking and PZT response measurements.

ELOG V3.1.3-